OpenStudy (anonymous):

please help me! A jewelry store is designing a gift box. the sum of the length, width and height is 12 inches. If the length is one inch greater in height, what should the dimensions of the box be to maximize its volume? what is the maximized volume?

OpenStudy (anonymous):

V = L W H = (H+1) W H = H^2W + WH L+W+H = 12 = (H+1) + W + H = W + 2H+1 W=11- 2H Make V(H) only and set dH/dH = 0 to get extreme's value of H check that (d/dH) (dV/dH) < 0 so answer is max not min

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

To maximize a function, want to get it to depend on nly one variable Here, we start with L, W, H. It seems that L=H+1, so we can make that substitution. We then have H and W. Putting W in terms of H, we now have volume as a function of H, which you maximize by setting dV(H)/dH = 0. If you do not know this technique, the problem is not suitable for you.

OpenStudy (anonymous):

can you show work? i am a visual learner

OpenStudy (anonymous):

please @douglaswinslowcooper

OpenStudy (anonymous):

@douglaswinslowcooper I think you will have to write them all using the same variable.

OpenStudy (anonymous):

im so confused by this im doing a chapter review

OpenStudy (anonymous):

Yes, the idea is to boil the relationships down to one variable, such as H.. I must leave.

OpenStudy (anonymous):

uuuh ok? & then whatr

OpenStudy (anonymous):

I think this guy would be able to help you. @iPwnBunnies @hartnn @FibonacciChick666 Any of these should come here.

OpenStudy (fibonaccichick666):

so if we can start a new? can you draw me a picture to represent the box?

OpenStudy (anonymous):

its just on my worksheet

OpenStudy (fibonaccichick666):

i mean can you yourself draw a picture

OpenStudy (anonymous):

idk how lol id just draw a 3d cube

OpenStudy (fibonaccichick666):

is a 3d cube a good enough representation for the box?

OpenStudy (fibonaccichick666):

in your mind?

OpenStudy (anonymous):

i guess

OpenStudy (fibonaccichick666):

so draw it

OpenStudy (anonymous):

i did

OpenStudy (fibonaccichick666):

on here

OpenStudy (anonymous):